Complete the ordered pairs for the equation y = 8x + 16.

( ,0),(1/8, ),(0, ),(-1/8, )

Answer by checkley75(3666)

Y=8X+16
0=8X+16
8X=-16
X=-16/8
X=-2 ANSWER. (-2,0)
Y=1/8*8+16
Y=1+16
Y=17 ANSWER. (1/8,17)
y=8*0+16
y=16 answer. (0,16)
y=-1/8*8+16
y=-1+16
y=15 answer. (-1/8,15)